OpenStudy (anonymous):

6 sin 112,5 sin 22,5 = ..... teach me please

ganeshie8 (ganeshie8):

hint : 112.5 = 180 - 67.5 22.5 = 90 - 67.5

OpenStudy (anonymous):

but these is "6sin" how?

ganeshie8 (ganeshie8):

\(\large 6~ \sin (112.5) ~~ \sin (22.5) \)

ganeshie8 (ganeshie8):

\(\large 6~ \sin (180 - 67.5) ~~ \sin (90-67.5) \)

ganeshie8 (ganeshie8):

can u take it from here ? :)

OpenStudy (anonymous):

i'm confuse. in my book the formula is 2sinAsinB=cos(A-B)-cos(A+B). and i really confuse :(

ganeshie8 (ganeshie8):

oh yes, thats a very nice formula. lets use that :)

ganeshie8 (ganeshie8):

\(\large 6~ \sin (112.5) ~~ \sin (22.5) \) \(\large 3 \times 2 ~ \sin (112.5) ~~ \sin (22.5) \)

ganeshie8 (ganeshie8):

u can apply the formula now ?

ganeshie8 (ganeshie8):

\(\large 6~ \sin (112.5) ~~ \sin (22.5) \) \(\large 3 \times 2 ~ \sin (112.5) ~~ \sin (22.5) \) \(\large 3 \times [\cos (112.5 - 22.5) - \cos(112.5 + 22.5)] \)

OpenStudy (anonymous):

3 x (cos90-cos 135) 3 x (0-cos135)? 3x -cos 135?

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

whats cos(135) ?

OpenStudy (anonymous):

i dont know haha

ganeshie8 (ganeshie8):

-3 cos (135) -3 cos ( 90 + 45) use cos (A + B) formula :)

ganeshie8 (ganeshie8):

cos(A + B) = cosA cosB - sinA sin B

ganeshie8 (ganeshie8):

-3 cos ( 90 + 45) -3 [cos90 cos45 - sin90 sin 45] -3 [0 - 1/sqrt(2)] 3/sqrt(2)

OpenStudy (anonymous):

oke thak ou very much :)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

you're so helpful :)

OpenStudy (anonymous):

can i ask once more?

ganeshie8 (ganeshie8):

sure ask...

OpenStudy (anonymous):

If θ is in quadrant 1 and tan θ = 5/12. The value of tan 4θ/5 = ...

ganeshie8 (ganeshie8):

not use if there is any smart way, il try like below :- tan θ = 5/12 => θ = arctan(5/12) = 22.62 degrees http://www.wolframalpha.com/input/?i=arctan%285%2F12%29+degrees%29

ganeshie8 (ganeshie8):

tan(4 θ/5) = tan(4 * 22.62/5) = 0.327 http://www.wolframalpha.com/input/?i=tan%284+*+22.62%2F5%29

OpenStudy (anonymous):

thank ypu ::)